# Dynamically Adaptive Reasoning via LLM-Guided MCTS for Efficient and Context-Aware KGQA

### v3

@@ Abstract
Knowledge Graph Question Answering (KGQA) aims to interpret natural language queries and perform structured reasoning over knowledge graphs by leveraging their relational and semantic structures to retrieve accurate answers. Recent KGQA methods primarily follow either retrieve-then-reason paradigm, relying on GNNs or heuristic rules for static paths extraction, or dynamic path generation strategies that use large language models (LLMs) with prompting to jointly perform retrieval and reasoning. However, the former suffers from limited adaptability due to static path extraction and lack of contextual refinement, while the latter incurs high computational costs and struggles with accurate path evaluation due to reliance on fixed scoring functions and extensive LLM calls. To address these issues, this paper proposes Dynamically Adaptive MCTS-based Reasoning (DAMR), a novel framework that integrates symbolic search with adaptive path evaluation for efficient and context-aware KGQA. DAMR employs a Monte Carlo Tree Search (MCTS) backbone guided by an LLM-based planner, which selects top-$k$ relevant relations at each step to reduce search space. To improve path evaluation accuracy, we introduce a lightweight Transformer-based scorer that performs context-aware plausibility estimation by jointly encoding the question and relation sequence through cross-attention, enabling the model to capture fine-grained semantic shifts during multi-hop reasoning. Furthermore, to alleviate the scarcity of high-quality supervision, DAMR incorporates a dynamic pseudo-path refinement mechanism that periodically generates training signals from partial paths explored during search, allowing the scorer to continuously adapt to the evolving distribution of reasoning trajectories. Extensive experiments on multiple KGQA benchmarks show that DAMR significantly outperforms state-of-the-art methods.
Introduction
@@
Large Language Models (LLMs) have demonstrated impressive reasoning capabilities across diverse tasks, including mathematical problem solving (Pei et al. 2025; Didolkar et al. 2024), commonsense inference (Toroghi et al. 2024), and open-domain question answering (Zhao et al. 2023). Despite their generalization ability, LLMs often struggle in domain-specific scenarios due to the lack of grounded external knowledge, resulting in factual hallucinations and high inference costs (Huang et al. 2025b; Wang et al. 2024). To address these limitations, recent efforts have explored integrating domain knowledge into LLM reasoning. A promising direction to overcome these limitations is Knowledge Graph Question Answering (KGQA) (Dammu, Naidu, and Shah 2025; Saxena, Tripathi, and Talukdar 2020; Choi et al. 2023), which integrates symbolic relational structures into the reasoning process to provide factual grounding and structural interpretability. By combining the expressiveness of natural language with the precision of knowledge graphs, KGQA offers a scalable solution to improve factual consistency, reasoning transparency, and answer reliability (Liu et al. 2025; Yao et al. 2025).
@@
Existing KGQA approaches can be broadly categorized into two main paradigms based on how they construct reasoning paths: retrieve-then-reason methods and dynamic path generation strategies. The first category adopts a retrieve-then-reason paradigm, where candidate reasoning paths are extracted using either Graph Neural Networks (GNNs) (Ma et al. 2025a; Yao et al. 2025) or rule-based heuristics (Fang et al. 2024) prior to answer prediction. However, these methods lack adaptability, as GNNs fail to incorporate question-specific semantics at inference time, while heuristic rules are inherently inflexible to support dynamic reasoning refinement (Liu et al. 2025; Yao et al. 2025). In contrast, dynamic path generation strategies unify retrieval and reasoning by constructing reasoning paths dynamically during question processing. These methods either prompt LLMs to iteratively generate paths via in-context learning or Chain-of-Thought (CoT) prompting (Sui et al. 2024; Li et al. 2024), or employ guided search techniques such as Monte Carlo Tree Search (MCTS) to incrementally expand paths with the aid of a path scorer (Ma et al. 2025b; Shen et al. 2025). Despite their flexibility, these approaches incur substantial computational overhead due to repeated LLM invocation and exhibit limited evaluation accuracy, as static scorers fail to capture the evolving semantics of reasoning paths (Chang et al. 2024; Shen et al. 2025).
@@
This paper investigates the design of an adaptive KGQA framework to address the challenges of computational inefficiency and limited path evaluation accuracy in dynamic reasoning. However, developing such a framework presents several key challenges: (1) How to modularize reasoning to reduce LLM overuse during search? A major source of computational inefficiency in dynamic KGQA lies in repeatedly invoking LLMs for both relation retrieval and reasoning during multi-hop path construction (Shen et al. 2025; Long et al. 2025). While methods such as CoT and MCTS provide flexible exploration, they tightly couple LLMs with each decision step, resulting in high inference costs and limited scalability. The key challenge is to design a modular reasoning framework that utilizes LLMs efficiently, guiding the search process without requiring direct involvement in every reasoning step. (2) How to accurately evaluate evolving reasoning paths? As multi-hop reasoning paths are incrementally constructed, their semantics evolve with each newly added relation and contextual information. However, existing methods typically rely on static scoring functions or shallow similarity metrics, which fail to capture the nuanced semantic shifts that occur throughout the reasoning process (Xu et al. 2024; Sui et al. 2024). This raises a key challenge: how to design a path evaluation model that adaptively captures fine-grained semantic changes conditioned on both the question and the evolving relation sequence. (3) How to train a reliable path evaluation model with limited supervision? Accurate path ranking in KGQA hinges on a well-calibrated evaluation model. However, dynamic reasoning methods typically produce a large number of incomplete or irrelevant paths, with only a small subset corresponding to valid reasoning trajectories. This results in highly imbalanced and noisy supervision, especially for multi-hop questions where successful paths are extremely sparse. Although reinforcement learning has been explored to mitigate this issue (Ma et al. 2024; Zhai et al. 2024), it frequently suffers from sparse rewards and unstable optimization. Therefore, the key challenge is how to construct meaningful learning signals from limited or implicit supervision to enable adaptive training of the path scorer.
@@
To address the above challenges, we propose Dynamically Adaptive MCTS-based Reasoning (DAMR), an efficient and adaptive reasoning framework that integrates symbolic search with context-aware semantic modeling to enable accurate and LLM-efficient multi-hop reasoning for KGQA. DAMR is built on an MCTS backbone, where an LLM-based planner dynamically guides path expansion by proposing semantically relevant relations at each step, significantly reducing search space and improving answer identification efficiency. To enable accurate and context-sensitive path evaluation, we introduce a lightweight Transformer-based scorer that estimates path plausibility by jointly encoding the question and relation sequence via cross-attention, effectively capturing evolving semantics during multi-hop reasoning. To address supervision scarcity, DAMR incorporates a dynamic pseudo-path mechanism that continuously adapts the scorer during search. Partial paths sampled from MCTS rollouts are ranked by predicted plausibility and converted into pseudo-path supervision pairs, amplifying learning signals from promising trajectories while suppressing noise from suboptimal ones. Extensive experiments on benchmark KGQA datasets demonstrate that DAMR significantly outperforms state-of-the-art baselines. Our contributions are summarized as follows:
@@
KGQA aims to enhance reasoning capabilities by incorporating external knowledge graphs to answer natural language questions (Choi et al. 2023; Xu et al. 2025). Existing KGQA approaches can be broadly classified into two categories: retrieve-then-reason and dynamic path generation. The first category extracts candidate reasoning paths using Graph Neural Networks (GNNs)(Ma et al. 2025a; Yao et al. 2025) or rule-based heuristics(Fang et al. 2024), followed by LLM-based answer generation. While GNNs learn embeddings to identify relevant paths and rule-based methods apply predefined patterns (Zhao et al. 2023; Liu et al. 2025), these approaches lack the flexibility to adapt dynamically to question-specific context during inference. In contrast, dynamic path generation methods, such as CoT prompting (Sui et al. 2024; Li et al. 2024) and MCTS (Ma et al. 2025b; Shen et al. 2025), unify retrieval and reasoning for more flexible exploration. However, they suffer from high computational overhead due to repeated LLM calls, and static scorers often fail to adapt to evolving path semantics (Long et al. 2025; Shen et al. 2025). To address these challenges, we propose an adaptive framework that integrates symbolic search with a fine-tuned evaluation model, aiming to improve both computational efficiency and reasoning accuracy in KGQA.
@@
Adaptive and Self-Improving Reasoning Models. A promising approach to developing adaptive reasoning models is to frame the process within a reinforcement learning (RL) paradigm, where an agent learns a policy to navigate a state space. Early methods such as DeepPath (Xiong, Hoang, and Wang 2017) and MINERVA (Das et al. 2018) used RL to discover reasoning paths by rewarding the agent only when a correct answer is reached. However, this leads to the sparse rewards problem—positive feedback arrives only after long action sequences, resulting in weak learning signals and poor exploration efficiency (Zhai et al. 2024; Chang et al. 2023). To address this challenge, an alternative is self-training via pseudo-labeling, where the model learns from its own high-confidence predictions (Lee et al. 2013; Xie et al. 2020). While commonly used in semi-supervised learning, pseudo-labeling proves especially effective in reasoning tasks with limited supervision (Wang et al. 2022; Huang et al. 2025a). Instead of relying on sparse terminal rewards, we leverage intermediate search paths as dynamic pseudo-paths, offering dense and adaptive supervision. This facilitates continual refinement of the path evaluator to better capture the evolving semantics of reasoning.
Preliminary
Problem Formulation
@@
We define Knowledge Graph Question Answering (KGQA) as the task of answering a natural language question $q$ by reasoning over a knowledge graph $\mathcal{K}$. The knowledge graph is typically represented as a set of triples $\mathcal{K}=\{(e_{s},r,e_{o})\}\subseteq\mathcal{E}\times\mathcal{R}\times\mathcal{E}$, where $\mathcal{E}$ and $\mathcal{R}$ denote the sets of entities and binary relations. The goal of KGQA is to find a set of answers $\mathcal{A}_{q}\subseteq\{(e_{1},r_{1},e_{2}),(e_{2},r_{2},e_{3})\cdots\}$ for question $q$, such that a semantic reasoning path through the graph leads from a topic entity to the correct answer. Formally, this is often framed as mapping $q$ to an executable query program $p_{q}$, where $\text{LLM}(p_{q}|{\mathcal{K}})=\mathcal{A}_{q}$.
Monte Carlo Tree Search
Monte Carlo Tree Search (MCTS) (Kocsis and Szepesvári 2006) is a heuristic search algorithm designed for optimal decision-making in structured search spaces. It incrementally builds a search tree through stochastic sampling and consists of four key stages:
Selection. Starting from the root, recursively select child nodes with the highest value according to the Upper Confidence Bound for Trees (UCT) criterion:
@@
where $w_{i}$ is the accumulated reward of node $i$, $n_{i}$ is the visit count of node $i$, NN is the visit count of its parent, and $C$ balances exploration and exploitation.
Expansion. Upon reaching a non-terminal leaf node, add one or more unexplored child nodes to the tree.
Simulation. From the newly added node, perform a random rollout (i.e., simulated trajectory) to a terminal state.
Backpropagation. Propagate the simulation outcome back up the tree, updating the statistics (e.g., visit count and reward) of each node along the path.
This iterative process incrementally refines the search tree, guiding the exploration toward high-reward paths.
Methodology
Overview of Framework
In this paper, we propose a dynamically adaptive reasoning framework DAMR for KGQA, as shown in Fig. 1. DAMR comprises three components: (1) LLM Guided Expansion. DAMR employs MCTS to incrementally expand reasoning paths, guided by an LLM-based planner that proposes relevant relations. This significantly reduces computational overhead and enhances efficiency in knowledge graph exploration; (2) Context-Aware Path Evaluation. To capture the evolving semantics of reasoning paths, DAMR employs a lightweight Transformer-based scorer with cross-attention to jointly encode the question and path embeddings. This enables context-sensitive evaluation and enhances the accuracy and relevance of multi-hop reasoning; (3) Path-based Dynamic Refinement. DAMR uses intermediate paths from MCTS as dynamic pseudo-paths to iteratively fine-tune the path evaluator, enhancing its ability to capture question-specific semantics and improving reasoning accuracy.
LLM Guided Expansion
A key challenge in KGQA is efficiently exploring the vast search space of multi-hop reasoning paths, especially under weak or no supervision. Existing methods often struggle to balance search efficiency and semantic relevance, resulting in either redundant exploration or missed correct paths. To address this, the LLM-Guided Expansion module employs MCTS (Kocsis and Szepesvári 2006) as the backbone for symbolic path expansion. At each step, an LLM proposes semantically relevant relations, narrowing the search space and improving path quality, while MCTS ensures a balanced trade-off between exploration and exploitation.
Specifically, each node in the MCTS represents a reasoning state anchored at a specific entity in the KG. Given the current state, possible actions correspond to selecting an outgoing relation to extend the reasoning path. During the Selection phase, nodes are scored using the UCT in Eq. (1), guiding the search to balance exploration and exploitation.
In the Expansion phase, we employ an LLM guided strategy to prioritize semantically meaningful path extensions. Given a specific entity $e_{i}$ in KG, we retrieve its associated outgoing relations $\mathcal{R}_{e_{i}}=\{r_{1},r_{2},\dots,r_{n}\}$. To focus the search on meaningful directions, we prompt an LLM with the question $q$ and the candidate relations $\mathcal{R}{e_{i}}$, selecting the top-$k$ relations most aligned with the question:
@@
These selected relations are then used to expand the current node. This LLM-guided expansion significantly reduces unnecessary branching and ensures that the search remains semantically focused and computationally efficient.
Context-Aware Path Evaluation

### v4

# DAMR: Efficient and Adaptive Context-Aware Knowledge Graph Question Answering with LLM-Guided MCTS

@@ Abstract
Knowledge Graph Question Answering (KGQA) aims to interpret natural language queries and perform structured reasoning over knowledge graphs by leveraging their relational and semantic structures to retrieve accurate answers. Existing methods primarily follow either the retrieve-then-reason paradigm, which relies on Graph Neural Networks (GNNs) or heuristic rules to extract static candidate paths, or dynamic path generation strategies that employ large language models (LLMs) with prompting to jointly perform retrieval and reasoning. However, the former lacks adaptability due to static path extraction and the absence of contextual refinement, while the latter suffers from high computational costs and limited evaluation accuracy because of their dependence on fixed scoring functions and repeated LLM calls. To address these issues, this paper proposes Dynamically Adaptive MCTS-based Reasoning (DAMR), a novel framework that integrates LLM-guided Monte Carlo Tree Search (MCTS) with adaptive path evaluation to enable efficient and context-aware KGQA. DAMR leverages MCTS as a backbone, where an LLM-based planner selects the top-$k$ semantically relevant relations at each expansion step to effectively reduce the search space. To enhance evaluation accuracy, we introduce a lightweight Transformer-based scorer that performs context-aware plausibility estimation by jointly encoding the question and relation sequence through cross-attention, thereby capturing fine-grained semantic shifts during multi-hop reasoning. Furthermore, to mitigate the scarcity of high-quality supervision, DAMR incorporates a dynamic pseudo-path refinement mechanism that periodically generates training signals from partial paths explored during search, enabling the scorer to continually adapt to the evolving distribution of reasoning trajectories. Extensive experiments on multiple KGQA benchmarks show that DAMR significantly outperforms state-of-the-art methods.
1 Introduction
@@
Large Language Models (LLMs) have demonstrated impressive reasoning capabilities across diverse tasks, including mathematical problem solving (pei2025mathfusion, ; didolkar2024metacognitive, ), commonsense inference (toroghi2024verifiable, ), and open-domain question answering (zhao2023divknowqa, ). Despite their generalization ability, LLMs often struggle in domain-specific scenarios due to the lack of grounded external knowledge, resulting in factual hallucinations and high inference costs (huang2025survey, ; wang2024factuality, ). To address these limitations, recent efforts have explored integrating domain knowledge into LLM reasoning. A promising direction to overcome these limitations is Knowledge Graph Question Answering (KGQA) (dammu2025dynamic, ; saxena2020improving, ; choi2023nutrea, ), which integrates relational structures into the reasoning process to provide factual grounding and structural interpretability. By combining the expressiveness of natural language with the precision of knowledge graphs, KGQA offers a scalable solution to improve factual consistency, reasoning transparency, and answer reliability (liu2025dual, ; yao2025learning, ).
@@
Existing KGQA approaches can be broadly categorized into two paradigms according to how reasoning paths are constructed: retrieve-then-reason methods and dynamic path generation strategies. In the former, candidate paths are extracted prior to answer prediction, typically using Graph Neural Networks (GNNs) (ma2025debate, ; yao2025learning, ) or rule-based heuristics (fang2024karpa, ). However, such methods exhibit limited adaptability, as GNNs fail to incorporate question-specific semantics during inference, while heuristic rules remain inherently inflexible and cannot support dynamic refinement of reasoning (liu2025dual, ; yao2025learning, ). In contrast, dynamic path generation strategies unify retrieval and reasoning by constructing paths dynamically during question processing. These methods either leverage LLMs to iteratively generate paths through in-context learning or Chain-of-Thought (CoT) (sui2024fidelis, ; li2024framework, ), or employ guided search techniques such as MCTS, where paths are incrementally expanded with a path scorer (ma2025deliberation, ; shen2025reasoning, ). While offering greater flexibility, such approaches suffer from substantial computational overhead due to repeated LLM calls and limited evaluation accuracy, as static scorers cannot capture the evolving semantics of reasoning paths (chang2024survey, ; lee2024zero, ).
@@
This paper investigates the design of an adaptive KGQA framework to address challenges of computational inefficiency and limited path evaluation accuracy in dynamic reasoning. However, developing such a framework raises three key challenges: (1) How to modularize reasoning to reduce LLM overuse in search? Inefficiency in dynamic KGQA stems from repeatedly invoking LLMs for relation retrieval and reasoning in multi-hop path construction (shen2025reasoning, ; long2025enhancing, ). Although methods such as CoT and MCTS enable flexible exploration, they bind LLMs to every decision step, causing high inference cost and poor scalability. The challenge is to design a modular framework that leverages LLMs efficiently, guiding search without direct involvement at each step. (2) How to accurately evaluate evolving reasoning paths? As multi-hop paths are incrementally constructed, their semantics evolve with each added relation and context. Yet existing methods rely on static scoring or shallow similarity metrics that fail to capture these semantic shifts (xu2024llm, ; sui2024fidelis, ). This highlights the challenge of designing a path evaluation model that adaptively captures fine-grained changes conditioned on both the question and evolving relation sequence. (3) How to train a reliable evaluation model with limited supervision? Accurate path ranking requires a well-calibrated scorer, yet dynamic reasoning generates many incomplete paths with only a few valid ones. This results in imbalanced, noisy supervision, especially for multi-hop questions where successful trajectories are scarce. While reinforcement learning has been explored (ma2024coevolving, ; zhai2024fine, ), it suffers from sparse rewards and unstable optimization. The key challenge is to construct reliable learning signals from limited supervision to support adaptive scorer training.
@@
To address these challenges, we propose Dynamically Adaptive MCTS-based Reasoning (DAMR), an efficient framework that integrates LLM-guided MCTS with context-aware semantic modeling for accurate and efficient KGQA. DAMR employs an MCTS backbone, where an LLM-based planner dynamically guides path expansion by selecting semantically relevant relations at each step, thereby reducing the search space and improving answer identification. For path evaluation, we introduce a lightweight Transformer-based scorer that jointly encodes the question and relation sequence via cross-attention, enabling context-sensitive plausibility estimation and capturing evolving semantics during multi-hop reasoning. To mitigate supervision scarcity, DAMR further incorporates a dynamic pseudo-path mechanism that continuously adapts the scorer during search: partial paths from MCTS rollouts are ranked by predicted plausibility and converted into pseudo-path supervision pairs, amplifying signals from promising trajectories while suppressing noise from suboptimal ones. Our contributions are summarized as follows:
@@ 2 Related work
KGQA aims to enhance reasoning capabilities by incorporating external knowledge graphs to answer natural language questions (choi2023nutrea, ; xu2025memory, ). Existing KGQA approaches can be broadly classified into two categories: retrieve-then-reason and dynamic path generation. The first category extracts candidate reasoning paths using Graph Neural Networks (GNNs)(ma2025debate, ; yao2025learning, ) or rule-based heuristics(fang2024karpa, ), followed by LLM-based answer generation. While GNNs learn embeddings to identify relevant paths and rule-based methods apply predefined patterns (zhao2023divknowqa, ; liu2025dual, ), these approaches lack the flexibility to adapt dynamically to question-specific context during inference. In contrast, dynamic path generation methods, such as CoT prompting (sui2024fidelis, ; li2024framework, ) and MCTS (ma2025deliberation, ; shen2025reasoning, ), unify retrieval and reasoning for more flexible exploration. However, they suffer from high computational overhead due to repeated LLM calls, and static scorers often fail to adapt to evolving path semantics (long2025enhancing, ; luo2025kbqa, ; ma2024think, ). To address these challenges, we propose an adaptive framework that integrates symbolic search with a fine-tuned evaluation model, aiming to improve both computational efficiency and reasoning accuracy in KGQA.
@@
Adaptive and Self-Improving Reasoning Models. A promising approach to developing adaptive reasoning models is to frame the process within a reinforcement learning (RL) paradigm, where an agent learns a policy to navigate a state space. Early methods such as DeepPath (xiong2017deeppath, ) and MINERVA (das2018go, ) used RL to discover reasoning paths by rewarding the agent only when a correct answer is reached. However, this leads to the sparse rewards problem, as positive feedback arrives only after long action sequences, resulting in weak learning signals and poor exploration efficiency (zhai2024fine, ; chang2023learning, ). To address this challenge, an alternative is self-training via pseudo-labeling, where the model learns from its own high-confidence predictions (lee2013pseudo, ; xie2020self, ). While commonly used in semi-supervised learning, pseudo-labeling proves especially effective in reasoning tasks with limited supervision (wang2022semi, ; huang2025mllm, ). Instead of relying on sparse terminal rewards, we leverage intermediate search paths as dynamic pseudo-paths, offering dense and adaptive supervision. This facilitates continual refinement of the path evaluator to better capture the evolving semantics of reasoning.
3 Methodology
3.1 Problem Formulation
@@
We define Knowledge Graph Question Answering (KGQA) as the task of answering a natural language question by reasoning over a knowledge graph (KG). The KG is typically represented as a set of triples $\mathcal{K}=\{(e_{s},r,e_{o})\}\subseteq\mathcal{E}\times\mathcal{R}\times\mathcal{E}$, where $\mathcal{E}$ and $\mathcal{R}$ denote the sets of entities and binary relations. The goal of KGQA is to find a set of answers $\mathcal{A}_{q}\subseteq\{(e_{1},r_{1},e_{2}),(e_{2},r_{2},e_{3})\cdots\}$ for question $q$, such that a reasoning path through the KG leads from a topic entity to the correct answer. Formally, this is often framed as mapping $q$ to an executable query program $p_{q}$, where $\text{LLM}(p_{q}|{\mathcal{K}})=\mathcal{A}_{q}$.
3.2 Overview of Framework
In this paper, we propose a dynamically adaptive reasoning framework DAMR for KGQA, as shown in Fig. 1. DAMR comprises three components: (1) LLM Guided Expansion. DAMR employs MCTS to incrementally expand reasoning paths, guided by an LLM-based planner that proposes relevant relations. This reduces computational overhead and enhances efficiency in KG exploration; (2) Context-Aware Path Evaluation. To capture the evolving semantics of reasoning paths, DAMR employs a lightweight Transformer-based scorer with cross-attention to jointly encode the question and path embeddings. This enables context-sensitive evaluation and improves the accuracy and relevance of multi-hop reasoning; (3) Path-based Dynamic Refinement. DAMR uses intermediate paths from MCTS as dynamic pseudo-paths to iteratively fine-tune the path evaluator, enhancing its ability to capture question-specific semantics and improving reasoning accuracy.
3.3 LLM Guided Expansion
A key challenge in KGQA is efficiently exploring the vast search space of multi-hop reasoning paths, especially under weak or no supervision. Existing methods often struggle to balance search efficiency and semantic relevance, resulting in either redundant exploration or missed correct paths (long2025enhancing, ; li2024simple, ). To address this, the LLM-Guided Expansion module employs MCTS (kocsis2006bandit, ) as the backbone for symbolic path expansion. At each step, an LLM proposes semantically relevant relations, narrowing the search space and improving path quality, while MCTS ensures a balanced trade-off between exploration and exploitation.
Specifically, each node in the MCTS represents a reasoning state anchored at a specific entity in the KG. Given the current state, possible actions correspond to selecting an outgoing relation to extend the reasoning path. During the Selection phase, the search traverses from the root to a leaf node by recursively choosing children with the highest Upper Confidence Bound for Trees (UCT) score (kocsis2006bandit, ), which is defined as:
@@ 3.3 LLM Guided Expansion
where $w_{i}$ denotes the accumulated reward of node $i$, $n_{i}$ its visit count, $N$ the visit count of its parent, and $C$ a constant that balances exploration and exploitation. This criterion guides the search to trade off between exploring new relations and exploiting promising paths. In the Expansion phase, the selected leaf node is expanded by retrieving the outgoing relations of its entity $e_{i}$, denoted as $\mathcal{R}_{e_{i}}=\{r_{1},r_{2},\dots,r_{n}\}$. To avoid exhaustive branching, we prompt an LLM with the question $q$ and candidate relations $\mathcal{R}_{e_{i}}$, selecting the top-$k$ relations most semantically aligned with $q$:
@@
These top-$k$ relations are then used to generate new child nodes, ensuring that expansion favors semantically meaningful directions and remains computationally efficient.
3.4 Context-Aware Path Evaluation